# SeniorTalk: A Chinese Conversation Dataset with Rich Annotations for Super-Aged Seniors

### v1

@@ Abstract
While voice technologies increasingly serve aging populations, current systems exhibit significant performance gaps due to inadequate training data capturing elderly-specific vocal characteristics like presbyphonia and dialectal variations. The limited data available on super-aged individuals in existing elderly speech datasets, coupled with overly simple recording styles and annotation dimensions, exacerbates this issue. To address the critical scarcity of speech data from individuals aged 75 and above, we introduce SeniorTalk, a carefully annotated Chinese spoken dialogue dataset. This dataset contains 55.53 hours of speech from 101 natural conversations involving 202 participants, ensuring a strategic balance across gender, region, and age. Through detailed annotation across multiple dimensions, it can support a wide range of speech tasks. We perform extensive experiments on speaker verification, speaker diarization, speech recognition, and speech editing tasks, offering crucial insights for the development of speech technologies targeting this age group.
SeniorTalk: A Chinese Conversation Dataset with Rich Annotations for Super-Aged Seniors
Yang Chen1+, Hui Wang1+, Shiyao Wang1, Junyang Chen1,Jiabei He1, Jiaming Zhou1, Xi Yang2, Yequan Wang2, Yonghua Lin2, Yong Qin1thanks: Yong Qin is the corresponding author. 1College of Computer Science, Nankai University, 2Beijing Academy of Artificial Intelligence, Beijing, China , Correspondence: 2120230617@mail.nankai.edu.cn, qinyong@nankai.edu.cn
First Author and Second Author contribute equally to this work.
@@ 1 Introduction
Dataset Age Style Annotation Features
Speaker Region Transcript Timestamp Accent Level Sound Event
ASR0060 55+ Reading North/South/Other $\checkmark$ N/A $\times$ $\times$
MECSD 55+ Reading $\times$ $\checkmark$ N/A $\times$ $\checkmark$
SeniorTalk 75+ Conversation Provincial $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
The rapid global aging population presents both significant challenges and opportunities in the development of technologies specifically designed for older adults, particularly those aged 75 and over (Scuteri and Nilsson, 2024; Chumuang et al., 2024). As the number of people in this ultra-high-age group continues to grow, it becomes increasingly important to enhance the accessibility and inclusion of speech-based technologies for this demographic (Young and Mihailidis, 2010; Portet et al., 2013). However, many state-of-the-art speech systems struggle to perform effectively within the elderly population, exhibiting biases on elderly vocal patterns. Kulkarni et al. (2024); Feng et al. (2021). For instance, existing speech recognition systems often show poor performance with older users, influenced by factors such as speech deterioration, hearing loss, health issues (Fraser et al., 2015), and the diversity of speech patterns among older adults (Geng et al., 2022; Hu et al., 2024). A key underlying reason for this challenge is the lack of datasets specifically addressing the unique needs of ultra-high-age individuals Hu et al. (2024), which hampers the development of robust foundation models and the application of tailored solutions.
@@
Although existing studies have made some efforts to collect geriatric speech data (Sekerina et al., 2024; Fukuda et al., 2020), significant limitations remain. First, current speech corpora predominantly focus on younger adults or healthy senior populations, with samples that consist mainly of scripted speech and exhibit standardized accents. For example, both the AISHELL-ASR0060 and MECSD datasets (Wang et al., 2019) adopt an unusually low age threshold of 55 years for defining elderly participants. Similarly, the S-JNAS corpus of elderly Japanese speech reports a mean speaker age of 67.6 years (Fukuda et al., 2020). These fall below the World Health Organization’s (WHO) geriatric classification designating 75+ years as the late elderly phase which is a period associated with progressive age-related decline in physiological function (Orimo et al., 2006).
@@
Second, the current collection paradigms and annotation methods of existing datasets further limit their practical applicability. A large portion of speech resources for the elderly primarily focuses on reading style (Wang et al., 2019; Iribe et al., 2015), which do not reflect the everyday communication scenarios that elderly people encounter in real life. Furthermore, many corpora are tailored for narrowly defined tasks such as automatic speech recognition (ASR) or pathology detection (Wang et al., 2019), thereby preventing comprehensive characterization of age-related vocal variations. These datasets also lack key features such as speaker diarization or dialect labels, which restrict their ability to support a wider array of use cases and hinder their robustness in addressing the diverse challenges of speech processing.
@@
To address these limitations, we introduce SeniorTalk, a Mandarin dataset consisting of spontaneous conversations among individuals aged 75 and older. As shown in Table 1, this dataset comprises natural conversational recordings from 202 native Chinese speakers, representing a rich diversity of regional, age, and gender demographics, and captured in authentic, real-world interaction settings. It effectively addresses the current gaps and limitations in datasets focused on elderly populations, particularly the underrepresentation of super-aged seniors and the lack of diversity in recording styles and annotation dimensions. Moreover, we conduct extensive experiments across various speech tasks, providing a benchmark specifically for the elderly population. By open-sourcing this corpus along with fine-grained metadata, we aim to bridge the vocal age gap and promote the development of equitable voice technologies for aging societies. SeniorTalk makes three key contributions to the field:
@@ 2 Related Work
BraPoRus (Sekerina et al., 2024) Brazilian Portuguese-Russian 59-98 Monologue, interview, … 1,500 170 2024 N
EARS (Fukuda et al., 2023) Japanese 70-99 Reading 123 13.4 2023 N
Improving S-JNAS (Fukuda et al., 2020) Japanese 65-99 Reading 221 31.7 2020 Y
@@
MECSD (Wang et al., 2019) Mandarin 55-85 Reading 85 110 2019 P
elderLUCID (Hazan et al., 2017) English 19-84 Reading 83 N 2017 N
Develop S-JNAS (Iribe et al., 2015) Japanese 60-98 Reading 100 9.2 2015 Y
ERES38 (Aman et al., 2013) French 68-98 Interview 22 17 2013 N
AD80 (Aman et al., 2013) French 62-94 - 43 4.7 2013 N
CCC (Pope and Davis, 2011) English 65+ Interview 600+ 800+ 2011 Y
@@
Several corpora have been developed to address the specific acoustic characteristics of elderly speech. Early work in this area includes the Japanese Newspaper Article Sentences Read Speech Corpus of the Aged (S-JNAS), a foundational resource for Japanese elderly speech research. Subsequent research has broadened the scope of investigation, with efforts such as the Carolinas Conversations Collection (CCC) (Pope and Davis, 2011) focusing on multiethnic elderly speakers with chronic conditions, offering valuable insights into how sociolinguistic factors and health status influence speech production. The AD80 and ERES38 corpora (Aman et al., 2013) advance French elderly speech analysis through distress detection benchmarks for ambient assisted living systems. Further corpus development has continued with the creation of specialized datasets, including a corpus of 100 elderly Japanese speakers designed to enhance human-robot interaction in elder care (Iribe et al., 2015). The elderLUCID project (Hazan et al., 2017) examines the complexities of speech communication in older adults, considering the interplay of hearing loss, phonation, articulation, and cognitive factors.
More recently, researchers have focused on specific languages and conditions, as exemplified by the Mandarin Speech Database for Early Dementia Detection (MECSD) (Wang et al., 2019) and the AISHELL-ASR0060 databasehttps://www.aishelltech.com/General_Datasets for elderly Mandarin speech. Improvements to existing corpora, such as S-JNAS, have also been explored, including the creation of acoustic models specifically designed for "super-elderly" speakers (Fukuda et al., 2020). The Elderly Multimodal Interpersonal Conversation (E-MIC)https://ai4robot.github.io/emic-en/ dataset expands the scope of analysis by incorporating multimodal data, including video and audio, to study turn-taking in elderly conversations. Further work, such as EARS (Fukuda et al., 2023), has continued to refine acoustic modeling techniques for super-elderly Japanese speakers, while the BraPoRus corpus (Sekerina et al., 2024) highlights the importance of preserving heritage languages and the challenges of remote data collection during the COVID-19 pandemic.
@@ 3.2 Statistics
Figure 1: A stacked bar chart showing the population count by age and gender. The chart illustrates the distribution of male and female populations across different age groups.
Figure 2: A dual-layer pie chart representing the distribution of provinces across the North and South regions. The inner layer highlights the overall proportions of North and South, while the outer layer shows the specific distribution of provinces within each region.
@@
We recruit 202 participants aged between 75 and 85 years, ensuring a diverse representation across different segments of the ultra-elderly population, and obtain the necessary consent for their participation. Specific authorization details are provided in Appendix A.1. Figure 1 illustrates the gender distribution across different age groups. The dataset includes 69 male and 131 female speakers, with a higher proportion of females. This gender imbalance stems from the relative ease of recruiting female participants during the data collection process, likely due to the higher life expectancy of women in the targeted age range. Geographically, as shown in Figure 2, 94 participants are from northern China, while 108 are from southern China, covering regions such as Beijing, Shanghai, and Sichuan. This regional diversity enhances dialect recognition models by exposing them to a wide range of linguistic patterns.
Figure 3: This histogram visualizes the frequency distribution of the given dataset, showing how the values are spread across different ranges.
@@
The dataset comprises spontaneous speech dialogues that span a wide range of topics, aimed at capturing the natural flow and diversity of real-life conversations. These dialogues are recorded using various mobile devices, with a distribution of 70% from Android devices and 30% from iOS smartphones. The topics of these dialogues are specifically chosen to address relevant issues for older adults, such as health, pets, retirement, and other related matters. Each conversation typically covers between one and three topics, with the distribution of topics visualized in Appendix A.2. Figure 3 illustrates the distribution of session durations, which range from 25 to 50 minutes, with the majority of sessions clustering around the 35-minute mark.
@@ 4 Experiments
Model # Params Dim Dev (%) PLDA Cosine similarity
EER (%) minDCF EER (%) minDCF
x-vector 4.2M 512 12.04 14.63 0.9768 19.26 0.9598
ResNet-TDNN 15.5M 256 4.372 10.88 0.8450 11.50 0.9196
ECAPA-TDNN 20.8M 192 8.86 11.54 0.9214 10.24 0.9582
@@ 4.1.1 Metrics
We adopt two scoring approaches: probabilistic linear discriminant analysis (PLDA) (Prince and Elder, 2007) and cosine similarity, with evaluation based on two metrics: (1) Equal Error Rate (EER): We define a threshold $\tau$ where the miss probability equals the false alarm probability. Specifically, if the similarity score is above this threshold, the system accepts that the speakers are the same person; if it is below this threshold, the system rejects the claim. This threshold is selected when the false acceptance rate equals the false rejection rate. (2) Minimum Detection Cost (minDCF): A cost-sensitive metric for evaluating speaker verification systems under application-specific conditions.
@@ 4.1.2 Baselines
We adapt three state-of-the-art speaker embedding systems pre-trained on VoxCeleb (Nagrani et al., 2017) through domain-specific fine-tuning via SpeechBrain (Ravanelli et al., 2021): x-vector architecture (Snyder et al., 2018)https://huggingface.co/speechbrain/spkrec-xvect-voxceleb, ECAPA-TDNN (Desplanques et al., 2020)https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, and ResNet-TDNN (Villalba et al., 2020)https://huggingface.co/speechbrain/spkrec-resnet-voxceleb. Detailed hyperparameters are listed in Appendix B.2.3.
@@ 4.1.3 Results and Analysis
Table 3 demonstrates two critical findings from our geriatric voice analysis: First, the strong baseline performance confirms our dataset’s suitability for aging voice biometrics. However, age-related vocal degradation (e.g., pitch instability, articulatory imprecision) introduces distinct challenges compared to pediatric voices, potentially affecting gender differentiation and speaker discriminability.
Second, the ECAPA-TDNN’s relative underperformance versus parameter-efficient model ResNet using PLDA testing method suggests susceptibility to overfitting in limited elderly speech data scenarios. This emphasizes the need for data augmentation strategies, regularization techniques and hyperparameter finetuning when deploying deep speaker models in geriatric voice applications.
@@ 4.2.2 Baselines
For the speaker diarization task, we use the PyAnnote toolkit (Bredin, 2023; Plaquet and Bredin, 2023) https://github.com/pyannote/pyannote-audio. This speaker diarization pipeline consists of three primary components: Voice Activity Detection (VAD), the speaker extractor, and clustering methods (e.g., K-Nearest Neighbors).
  • VAD: We employ the pyannote/segmentation-3.0 model https://huggingface.co/pyannote/segmentation-3.0, an end-to-end neural architecture for joint speech activity detection and speaker segmentation. Since false alarm and missed detection rates remain consistent across experimental conditions due to fixed VAD parameters, Table 4 exclusively reports confusion errors and DER metrics.
  • Speaker extractor: For the speaker extractor module, we replace the default ResNet-34-LM extractor from PyAnnote https://huggingface.co/pyannote/wespeaker-voxceleb-resnet34-LM with fine-tuned versions of x-vector, ECAPA-TDNN, and ResNet-TDNN architectures, all adapted for speaker verification. This yields four distinct experimental configurations, as detailed in Table 4.

### v2

@@ Abstract
While voice technologies increasingly serve aging populations, current systems exhibit significant performance gaps due to inadequate training data capturing elderly-specific vocal characteristics like presbyphonia and dialectal variations. The limited data available on super-aged individuals in existing elderly speech datasets, coupled with overly simple recording styles and annotation dimensions, exacerbates this issue. To address the critical scarcity of speech data from individuals aged 75 and above, we introduce SeniorTalk, a carefully annotated Chinese spoken dialogue dataset. This dataset contains 55.53 hours of speech from 101 natural conversations involving 202 participants, ensuring a strategic balance across gender, region, and age. Through detailed annotation across multiple dimensions, it can support a wide range of speech tasks. We perform extensive experiments on speaker verification, speaker diarization, speech recognition, and speech editing tasks, offering crucial insights for the development of speech technologies targeting this age group. Code is available at https://github.com/flageval-baai/SeniorTalk and data at https://huggingface.co/datasets/evan0617/seniortalk.
@@ 1 Introduction
The rapid global aging population presents both significant challenges and opportunities in the development of technologies specifically designed for older adults, particularly those aged 75 and over (chumuang2024voice; SCUTERI202435). As the number of people in this ultra-high-age group continues to grow, it becomes increasingly important to enhance the accessibility and inclusion of speech-based technologies for this demographic (portet2013design; young2010difficulties). However, many state-of-the-art speech systems struggle to perform effectively within the elderly population, exhibiting biases on elderly vocal patterns. kulkarni-etal-2024-balancing; feng2021quantifying. For instance, existing speech recognition systems often show poor performance with older users, influenced by factors such as speech deterioration, hearing loss, health issues (fraser2015linguistic), and the diversity of speech patterns among older adults (geng2022speaker; 10584335). A key underlying reason for this challenge is the lack of datasets specifically addressing the unique needs of ultra-high-age individuals 10584335, which hampers the development of robust foundation models and the application of tailored solutions.
@@
Although existing studies have made some efforts to collect geriatric speech data (sekerina2024brazilian; fukuda-etal-2020-improving), significant limitations remain. First, current speech corpora predominantly focus on younger adults or healthy senior populations, with samples that consist mainly of scripted speech and exhibit standardized accents. For example, both the AISHELL-ASR0060 and MECSD datasets (wang19q_interspeech) adopt an unusually low age threshold of 55 years for defining elderly participants. In particular, these data sets show extremely low proportions of participants aged 75 to 85 years: only 0.8% for AISHELL-ASR0060 and 9.4% for MECSD, which means that the oldest population (75+ years) is severely underrepresented. Similarly, the S-JNAS corpus of elderly Japanese speech reports a mean speaker age of 67.6 years (fukuda-etal-2020-improving). These fall below the World Health Organization’s (WHO) geriatric classification designating 75+ years as the late elderly phase which is a period associated with progressive age-related decline in physiological function (orimo2006reviewing).
@@
Second, the current collection paradigms and annotation methods of existing datasets further limit their practical applicability. A large portion of speech resources for the elderly primarily focuses on reading style (wang19q_interspeech; 7357859), which do not reflect the everyday communication scenarios that elderly people encounter in real life. Furthermore, many corpora are tailored for narrowly defined tasks such as automatic speech recognition (ASR) or pathology detection (wang19q_interspeech), thereby preventing comprehensive characterization of age-related vocal variations. These datasets also lack key features such as speaker diarization or dialect labels, which restrict their ability to support a wider array of use cases and hinder their robustness in addressing the diverse challenges of speech processing.
@@
To address these limitations, we introduce SeniorTalk, the only freely available open-source Mandarin speech dataset consisting of spontaneous conversations among individuals aged 75 and older. As shown in Table 1, this dataset comprises natural conversational recordings from 202 native Chinese speakers, representing a rich diversity of regional, age, and gender demographics, and captured in authentic, real-world interaction settings. It effectively addresses the current gaps and limitations in datasets focused on elderly populations, particularly the underrepresentation of super-aged seniors and the lack of diversity in recording styles and annotation dimensions. Moreover, we conduct extensive experiments across various speech tasks, providing a benchmark specifically for the elderly population. By open-sourcing this corpus along with fine-grained metadata, we aim to bridge the vocal age gap and promote the development of equitable voice technologies for aging societies. SeniorTalk makes three key contributions to the field:
@@
Dataset Mean Age Style Annotation Features Age Group Distribution (%)
Region Transcript Timestamp Accent Sound 55-65 65-75 75-85
ASR0060 60.42 Reading North/South/Other $\checkmark$ N/A $\times$ $\times$ 56.7 42.5 0.8
MECSD 67.27 Reading $\times$ $\checkmark$ N/A $\times$ $\checkmark$ 17.6 72.9 9.4
SeniorTalk 79.5 Conversation Provincial $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0 0 100
@@ 2 Related Work
BraPoRus (sekerina2024brazilian) Brazilian Portuguese-Russian 59-98 Monologue, interview, … 1,500 170 2024 N
EARS (FUKUDA2023101424) Japanese 70-99 Reading 123 13.4 2023 N
Improving S-JNAS (fukuda-etal-2020-improving) Japanese 65-99 Reading 221 31.7 2020 Y
@@
MECSD (wang19q_interspeech) Mandarin 55-85 Reading 85 110 2019 P
elderLUCID (hazanelderlucid) English 19-84 Reading 83 N 2017 N
Develop S-JNAS (7357859) Japanese 60-98 Reading 100 9.2 2015 Y
ERES38 (aman2013speech) French 68-98 Interview 22 17 2013 N
AD80 (aman2013speech) French 62-94 - 43 4.7 2013 N
CCC (PopeDavis+2011+143+161) English 65+ Interview 600+ 800+ 2011 Y
@@
Several corpora have been developed to address the specific acoustic characteristics of elderly speech. Early work in this area includes the Japanese Newspaper Article Sentences Read Speech Corpus of the Aged (S-JNAS), a foundational resource for Japanese elderly speech research. Subsequent research has broadened the scope of investigation, with efforts such as the Carolinas Conversations Collection (CCC) (PopeDavis+2011+143+161) focusing on multiethnic elderly speakers with chronic conditions, offering valuable insights into how sociolinguistic factors and health status influence speech production. The AD80 and ERES38 corpora (aman2013speech) advance French elderly speech analysis through distress detection benchmarks for ambient assisted living systems. Further corpus development has continued with the creation of specialized datasets, including a corpus of 100 elderly Japanese speakers designed to enhance human-robot interaction in elder care (7357859). The elderLUCID project (hazanelderlucid) examines the complexities of speech communication in older adults, considering the interplay of hearing loss, phonation, articulation, and cognitive factors.
More recently, researchers have focused on specific languages and conditions, as exemplified by the Mandarin Speech Database for Early Dementia Detection (MECSD) (wang19q_interspeech) and the AISHELL-ASR0060 database [^1] https://www.aishelltech.com/General_Datasets for elderly Mandarin speech. Improvements to existing corpora, such as S-JNAS, have also been explored, including the creation of acoustic models specifically designed for "super-elderly" speakers (fukuda-etal-2020-improving). The Elderly Multimodal Interpersonal Conversation (E-MIC) [^2] https://ai4robot.github.io/emic-en/ dataset expands the scope of analysis by incorporating multimodal data, including video and audio, to study turn-taking in elderly conversations. Further work, such as EARS (FUKUDA2023101424), has continued to refine acoustic modeling techniques for super-elderly Japanese speakers, while the BraPoRus corpus (sekerina2024brazilian) highlights the importance of preserving heritage languages and the challenges of remote data collection during the COVID-19 pandemic.
@@ 3.1 Dataset overview
(a) Population by age and gender
(b) Duration distribution
Figure 1: Data analysis: (a) Age-gender structure; (b) Duration distribution.
@@ 3.2 Statistics
We recruit 202 participants aged between 75 and 85 years, ensuring a diverse representation across different segments of the ultra-elderly population, and obtain the necessary consent for their participation. Specific authorization details are provided in Appendix A.1. Figure 1a illustrates the gender distribution across different age groups. The dataset includes 67 male and 135 female speakers, with a higher proportion of females. This gender imbalance stems from the relative ease of recruiting female participants during the data collection process, likely due to the higher life expectancy of women in the targeted age range. Geographically, as shown in Figure 2, 94 participants are from northern China, while 108 are from southern China, covering regions such as Beijing, Shanghai, and Sichuan. This regional diversity enhances dialect recognition models by exposing them to a wide range of linguistic patterns.
Figure 2: This histogram visualizes the region distribution of the given dataset, showing how the values are spread across different region.
@@
The dataset comprises spontaneous speech dialogues that span a wide range of topics, aimed at capturing the natural flow and diversity of real-life conversations. These dialogues are recorded using various mobile devices, with a distribution of 70% from Android devices and 30% from iOS smartphones. The topics of these dialogues are specifically chosen to address relevant issues for older adults, such as health, pets, retirement, and other related matters. Each conversation typically covers between one and three topics, with the distribution of topics visualized in Appendix A.2. Figure 1b illustrates the distribution of session durations, which range from 25 to 50 minutes, with the majority of sessions clustering around the 35-minute mark.
@@ 4.1.1 Metrics
We adopt two scoring approaches: probabilistic linear discriminant analysis (PLDA) (prince2007probabilistic) and cosine similarity, with evaluation based on two metrics: (1) Equal Error Rate (EER): We define a threshold $\tau$ where the miss probability equals the false alarm probability. Specifically, if the similarity score is above this threshold, the system accepts that the speakers are the same person; if it is below this threshold, the system rejects the claim. This threshold is selected when the false acceptance rate equals the false rejection rate. (2) Minimum Detection Cost (minDCF): A cost-sensitive metric for evaluating speaker verification systems under application-specific conditions.
@@ 4.1.2 Baselines
We adapt three state-of-the-art speaker embedding systems pre-trained on VoxCeleb (nagrani17_interspeech) through domain-specific fine-tuning via SpeechBrain (speechbrain): x-vector architecture (xvector2018), [^3] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb ECAPA-TDNN (ecapa2020), [^4] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb and ResNet-TDNN (resnet2020). [^5] https://huggingface.co/speechbrain/spkrec-resnet-voxceleb Detailed hyperparameters are listed in Appendix B.2.3.
Model # Params Dim Dev (%) PLDA Cosine similarity
EER (%) minDCF EER (%) minDCF
x-vector 4.2M 512 12.04 17.53 0.9032 22.38 0.8964
ResNet-TDNN 15.5M 256 4.37 11.39 0.7259 13.15 0.8225
ECAPA-TDNN 20.8M 192 5.85 8.56 0.6263 5.66 0.6121
@@ 4.1.3 Results and Analysis
Table 3 demonstrates two critical findings from our geriatric voice analysis: First, the strong fintuned models’ performance relative to the pretrained model validates the suitability of our dataset for aging voice biometric tasks. This suggests that age-related vocal degradation (e.g., pitch instability, articulatory imprecision) introduces distinct challenges compared to pediatric voices, potentially affecting gender differentiation and speaker discriminability. The pretrained model results without fine-tuning are provided in Appendix B.3.2 for comparison.
Second, despite achieving the best performance on the test set, the ECAPA-TDNN underperforms the ResNet-TDNN on the development set, indicating potential overfitting risks for elderly speech data. This emphasizes the need for data augmentation strategies, regularization techniques and hyperparameter tuning when deploying deep speaker models in geriatric voice applications.
@@ 4.2.2 Baselines
For the speaker diarization task, we use the PyAnnote toolkit (Bredin23; Plaquet23). [^6] https://github.com/pyannote/pyannote-audio This speaker diarization pipeline consists of three primary components: Voice Activity Detection (VAD), the speaker extractor, and clustering methods (e.g., K-Nearest Neighbors).
  • VAD: We employ the pyannote/segmentation-3.0 model, [^7] https://huggingface.co/pyannote/segmentation-3.0 an end-to-end neural architecture for joint speech activity detection and speaker segmentation. Since false alarm and missed detection rates remain consistent across experimental conditions due to fixed VAD parameters, Table 4 exclusively reports confusion errors and DER metrics.
  • Speaker extractor: For the speaker extractor module, we replace the default ResNet-34-LM extractor from PyAnnote  [^8] https://huggingface.co/pyannote/wespeaker-voxceleb-resnet34-LM with fine-tuned versions of x-vector, ECAPA-TDNN, and ResNet-TDNN architectures, all adapted for speaker verification. This yields four distinct experimental configurations, as detailed in Table 4.